Assume that LM curve equation as( M/P)=d1y-d2i . solve for equilibrium output.derive the multiplier
1
Expert's answer
2021-07-26T16:20:02-0400

"( M/P)=d1y-d2i\\
(1-c_1-b_1)Y = c_0-c_1T+b_o-b_2 i +G\\
Multiplier = \frac{1}{1-c_1-b_1}"

Substituting for interest

"Y= \frac{1}{(1-c_1-b_1+b_2d_1)/d_2}[c_0-c_1T+b_0+ (\frac{b_2}{d_2})(\frac{M}{P})+G]"

Multiplier becomes

"\frac{1}{(1-c_1-b_1+b_2d_1)/d_2}"